### v2

# SHAP values via sparse Fourier representation

Abstract
SHAP (SHapley Additive exPlanations) values are a widely used method for local feature attribution in interpretable and explainable AI. We propose an efficient two-stage algorithm for computing SHAP values in both black-box setting and tree-based models. Motivated by spectral bias in real-world predictors, we first approximate models using compact Fourier representations, exactly for trees and approximately for black-box models. In the second stage, we introduce a closed-form formula for exactly computing SHAP values using the Fourier representation, that “linearizes” the computation into a simple summation and is amenable to parallelization. As the Fourier approximation is computed only once, our method enables amortized SHAP value computation, achieving significant speedups over existing methods and a tunable trade-off between efficiency and precision.
1 Introduction
Interpretability of machine learning models is paramount, especially in high-stakes applications in areas such as medicine, fraud detection, or credit scoring. This is crucial to the extent that in Europe, the General Data Protection Regulation (GDPR) mandates the legal right to an explanation of algorithmic decisions [Voigt and Von dem Bussche, 2017]. Say we are given a predictor/model $h:\mathcal{X}^{n}\rightarrow\mathbb{R}$ which maps an input (data) instance $x^{*}\in\mathcal{X}^{n}$ to a prediction $h(x^{*})$. Instance-wise a.k.a. local feature attribution methods assign “importances” to each of the features $x_{i}^{*}\in\mathcal{X}$ of the instance $x^{*}$ which quantify how influential that feature was in the model predicting the value $h(x^{*})$.
A widely used method for deriving attributions (importances) is the notion of SHapley Additive exPlanations, commonly referred to simply as SHAP values. Originally, the notion of Shapley values was introduced in the seminal work of Shapley [1952] in the context of cooperative game theory. The Shapley value is a mathematically well-founded and “fair” way of distributing a reward among all the members of a group playing a cooperative game and it is computed based on the rewards that would be received for all possible coalitions. The Shapley value is the unique way of distributing the reward that satisfies several reasonable mathematical properties that capture a notion of fairness [Shapley, 1952]. In the context of machine learning and statistics, the players become features, the reward is the prediction of the predictor $h$ and the SHAP value is the “contribution” or “influence” of that feature on the prediction. Shapley values are widely used due to their mathematical soundness and desirable properties Gromping [2007], Štrumbelj et al. [2009], Owen [2014], Datta et al. [2016], Owen and Prieur [2017], Lundberg and Lee [2017], Lundberg et al. [2020] Aas et al. [2021].
Despite their prevalence, computing SHAP values is challenging, as it involves an exponential summation, i.e. a summation over exponentially many terms , see Equation 1. This is because the formula accounts for a feature’s importance in the context of all possible “coalitions” of other features, and therefore the formula covers all possible subsets of other features. Therefore, approximating them and speeding up the computation has received attention in a variety of settings. SHAP value computation can easily dominate the computation time of industry-level machine learning solutions on datasets with millions or more entries [Yang, 2021]. Yang [2021] point out that industrial applications sometimes require hundreds of millions of samples to be explained. Examples include feed ranking, ads targeting, and subscription propensity models. In these modeling pipelines, spending tens of hours in model interpretation becomes a significant bottleneck [Yang, 2021] and one usually needs to resort to multiple cores and parallel computing.
Significant work has gone into speeding up the computation of SHAP values for a variety of settings. In the (ensemble of) trees setting, full access to the tree structure is assumed. Yang [2021], Bifet et al. [2022] provide theoretical and practical computational speedups to the classic TreeShap [Lundberg et al., 2020]. Similar to these results, in this work, we provide significant speedups for the tree setting over previous methods.
As opposed to the tree setting, which is a “white box” setting, in the model-agnostic a.k.a black-box setting, we only have query access to the model. Here, our only means of access to the predictor is that we can pick an arbitrary $x\in\mathcal{X}^{n}$ and query the predictor for its value $h(x)$. The usual approach here is to approximate the exponential sum of the SHAP value computations using stochastic sampling [Covert and Lee, 2020, Mitchell et al., 2022a, Lundberg and Lee, 2017]. In this setting, Covert and Lee [2020], Mitchell et al. [2022a] provide sampling methods that require fewer queries to the black-box compared to vanilla KernelShap [Lundberg and Lee, 2017] for equal approximation accuracy. FastShap Jethani et al. [2021], which introduces a method for estimating Shapley values in a single forward pass using an end-to-end learned explainer neural network model. Our algorithm FourierShap falls into the query-access black-box setting as well.
However, we take a different approach. We are guided by the key insight that many models used in practice have a “spectral bias”. Yang and Salman [2019], Valle-Perez et al. [2018] provably and experimentally show that fully connected neural networks with binary (zero-one) inputs learn low-degree – and therefore sparse – functions in a basis called the Walsh-Hadamard a.k.a Fourier basis. It is well known that the Walsh-Hadamard transform (WHT) of an ensemble of $T$ trees of depth $d$ is also of degree at most $d$ and moreover, $k=O(T4^{d})$-sparse [Kushilevitz and Mansour, 1993, Mansour, 1994].
Our contributions: Guided by the aforementioned insights, we provide an algorithm to compute SHAP values, using the Fourier representation of the tree or black-box model. We first approximate the black-box function by taking its sparse Fourier transform. We theoretically justify, and show through extensive experiments, that for many real-world models such as fully connected neural networks and (ensembles of) trees this representation is accurate. Subsequently, we prove that SHAP values for a single Fourier basis function admit a closed-form expression not involving an exponential summation. Therefore, using the Fourier representation we overcome the exponential sum and can utilize compute power effectively to compute SHAP values. Furthermore, the closed-form expression we derive effectively “linearizes” the computation into a simple summation and is amenable to parallelization on multiple cores or a GPU. The Fourier approximation step is only done once, therefore FourierShap amortizes the cost of computing explanations for many inputs. Subsequently, SHAP computations using the Fourier approximation are orders of magnitude faster compared to other black-box SHAP approximation methods such as KernelShap and other variations of it which all involve a computationally expensive optimization. We show speedups over other methods such as DeepLift [Shrikumar et al., 2017] and FastShap [Jethani et al., 2021] as well. In addition to the speedup, our algorithm enables a reliable continuous trade-off between computation and accuracy, controlled in a fine manner by the sparsity of the Fourier approximation.
2 Background
This section reviews the notion of SHAP values, and sparse and low-degree Fourier transforms.
2.1 Shapley values
In game theory, a cooperative game is a function $v:2^{[n]}\rightarrow\mathbb{R}$ that maps a subset (coalition) $S\subseteq[n]$ of a group of players $[n]=\{1,\dots,n\}$ to their total reward (when they cooperate). If all the players cooperate they win a total reward of value $v([n])$. The main question is how they would distribute this reward among themselves. Shapley [1952] resolved this problem by a deriving a unique value based on “fairness axioms” proposed in his seminal work [Shapley, 1952]. The Shapley value of player $i\in[n]$ is:
$\phi_{i}(v)=\frac{1}{n}\sum\limits_{S\subseteq[n]\setminus\{i\}}\frac{v(S\cup\{i\})-v(S)}{\binom{n-1}{|S|}}$ (1)
Intuitively, one can view the term $v(S\cup\{i\})-v(S)$ as the marginal contribution of player $i$ when they are added to the coalition $S$. This marginal value is weighted according to the number of permutations the leading $|S|$ players and trailing $n-|S|-1$ players can form.
In the machine learning context, we have a predictor $h:\mathcal{X}^{n}\rightarrow\mathbb{R}$ mapping an $n$-dimensional feature vector to a value. In this context, the players become features $x_{i}\in\mathcal{X}$ and the reward is the prediction of the predictor $h$ and the Shapley value is the “contribution” or “influence” of the $i$’th feature on the prediction. We define $v(S)$ accordingly to capture this notion [Lundberg and Lee, 2017]:
$v(S)=\mathbb{E}_{\bm{x_{[n]\setminus S}}\sim p(\bm{x_{[n]\setminus S}}))}[h(x^{*}_{S},\bm{x_{[n]\setminus S}})],$
where $x^{*}\in\mathcal{X}^{n}$ is the instance we are explaining. This definition implicitly captures the way we handle the missing features (feature not present in the coalition): we integrate the missing features with respect to the marginal distribution $p(\bm{x_{[n]\setminus S}}))$. In practice, the marginalization is performed with an empirical distribution by taking a subset of the training data as background dataset.
The choice of which distribution to average the missing features from has been discussed thoroughly in the relevant literature. As mentioned before, in this work, we focus on the SHAP values as defined in KernelShap introduced by Lundberg and Lee [2017], Lundberg et al. [2020] also known as “Interventional” [Janzing et al., 2020, Van den Broeck et al., 2021] or “Baseline” [Sundararajan and Najmi, 2020] SHAP, where the missing features are integrated out from the marginal distribution $p(\bm{x_{[n]\setminus S}})$, as opposed to the conditional distribution $p(\bm{x_{[n]\setminus S}}|\bm{x_{S}})$. We refer the reader to Appendix A.3 for a comprehensive overview of the literature discussing these two notions and their conceptual differences.
Since we will be using the well-known KernelShap [Lundberg and Lee, 2017] and its variant LinRegShap [Covert and Lee, 2020] as a baseline we briefly review their method here. Lundberg and Lee [2017] propose the “least squares characterization” of SHAP values. They prove that SHAP values are the solution to the following minimization problem:
$\beta_{0}^{*},\dots,\beta_{n}^{*}\triangleq\arg\min\limits_{\beta_{0},\dots,\beta_{n}}\sum\limits_{0<|S|<n}\frac{n-1}{{\binom{n}{|S|}}|S|(n-|S|)}\left(\beta_{0}+\sum\limits_{i\in S}\beta_{i}-v(S)\right)\\ s.t.:\;\;\;\beta_{0}=v(\{\}),\beta_{0}+\sum\limits_{i=1}^{n}\beta_{i}=v([n])$
Then $\phi_{i}(v)=\beta_{i}^{*}$.
Unfortunately, the above optimization still involves an exponential sum. Therefore, Lundberg and Lee [2017] propose to sample subsets $S$ uniformly at random. Covert and Lee [2020], Mitchell et al. [2022a] provide better ways of sampling and solving the optimization to get approximations with lower variances and biases. Nevertheless, all these methods require solving a least squares minimization subject to constraints for each explained instance $x^{*}$ and, therefore, are computationally expensive.
Later, we also compare our method to FastSHAP [Jethani et al., 2021], a model-agnostic algorithm for computing SHAP values. In FastSHAP, a parametric explainer $\phi$ (e.g., MLP) is trained to directly generate SHAP values given data samples. Since training is only done once, this, similar to us, amortizes the cost of generating SHAP values across multiple instances. Computing SHAP values only requires a forward pass on the trained model and therefore can be done very quickly.
2.2 Fourier representations
Here we review the notions of the Fourier basis and what we mean by sparsity. We will later use sparse Fourier representation of functions to compute SHAP values. In this work, we focus on the setting where the inputs to the black-box function (predictor) are binary, i.e., $\mathcal{X}^{n}=\{0,1\}^{n}$. This means, we assume we have binary features and/or categorical features, through standard one-hot representations  [^1] Continuous features can be discretized into quantiles, enabling their transformation into categorical features.. The Fourier representation of the pseudo-boolean function $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ is the unique expansion of $h$ as follows: $h(x)=\frac{1}{\sqrt{2^{n}}}\sum\limits_{f\in\{0,1\}^{n}}\widehat{h}(f)(-1)^{\langle f,x\rangle}$.
The inner product of two vectors $f,x\in\{0,1\}^{n}$ is defined as: $\langle f,x\rangle\equiv\sum_{i}^{n}f_{i}x_{i},\forall f,x\in\{0,1\}^{n}$.
The unique function $\widehat{h}:\{0,1\}^{n}\rightarrow\mathbb{R}$ is called the Fourier transform of $h$. For any $f\in\{0,1\}^{n}$, $\widehat{h}(f)$ is called the Fourier coefficient corresponding to the frequency $f$. The family of functions $\frac{1}{\sqrt{2^{n}}}\Psi_{f}(x)=(-1)^{\langle f,x\rangle},f\in\{0,1\}^{n}$ are the $2^{n}$-many Fourier basis functions. These basis functions are orthonormal: $\sum\limits_{x\in\{0,1\}^{n}}\Psi_{f}(x)\Psi_{f^{\prime}}(x)=\begin{cases}0&f\neq f^{\prime}\\ 1&f=f^{\prime}\end{cases},\;\;\;f,f^{\prime}\in\{0,1\}^{n}$. Therefore, they form a basis for the vector space of all pseudo-boolean functions $h:\{0,1\}^{n}\rightarrow\mathbb{R}$.
We define the support of $h$ to be $\text{supp}(h)=\{f\in\{0,1\}^{n}|\widehat{h}(f)\neq 0\}$. We say that a function $h$ is $k$-sparse if at most $k$ of the $2^{n}$ Fourier coefficients $\widehat{h}(f)$ are non-zero, i.e., $|\text{supp}(h)|\leq k$. The degree of a vector $f\in\{0,1\}^{n}$ is denoted by $\deg(f)$ and is defined as the number of ones in the vector. For example, if $n=5$ then $f=(1,0,0,1,0)$ is a vector of degree $\deg(f)=2$. We say a function is degree $d$ when the frequencies $f\in\{0,1\}^{n}$ corresponding to non-zero Fourier coefficients are of degree less or equal to $d$ i.e. $\forall f\in\text{supp}(h)$ it holds that $\text{deg}(f)\leq d$.
By definition of the Fourier basis, a $k$-sparse degree $d$ function can be written as a summation of $k$ (Fourier basis) functions, each one depending on at most $d$ input variables. The converse is also true:
Proposition 1.
Assume $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ can be decomposed as follows: $h(x)=\sum\limits_{i=1}^{p}h_{i}(x_{S_{i}}),S_{i}\subseteq[n]$. That is, each function $h_{i}:\{0,1\}^{|S_{i}|}\rightarrow\mathbb{R}$ depends on at most $|S_{i}|$ variables. Then, $h$ is $k=O(\sum\limits_{i=1}^{p}2^{|S_{i}|})$-sparse and of degree $d=\max(|S_{1}|,\dots,|S_{p}|)$. (Proof in Appendix B.1)
The sparsity $k$ and degree $d$ capture a notion of complexity for the underlying function. Intuitively speaking, the sparsity factor $k$ puts a limit on the number of functions in the decomposition, and the degree $d$ puts a limit on the order of interactions among the input variables.
One can see from Proposition 1, that modular functions, i.e., functions that can be written as a sum of functions each depending on exactly one variable, are $k=O(n)$-sparse and of degree $d=1$. A slightly more “complex” function capturing second-order interactions among the input variables, i.e., a function that can be written as a sum where each term depends on at most two variables is going to be $k=O(n^{2})$-sparse and of degree $d=2$. The following proposition generalizes this result.
Proposition 2.
Let, $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ be a pseudo-boolean function and let $d\in\mathbb{N}$ be some constant (w.r.t. $n$). If $h$ is of degree $d$, then, it is $k=O(n^{d})$-sparse. (Proof in Appendix B.1)
This proposition formally shows that limiting the order of interactions among the input variables implies an upper bound on the sparsity.
3 Many real-world black-box predictors have sparse Fourier transforms
In this section, we discuss the sparsity of the Fourier transforms of ensembles of trees and why neural networks can be approximated by a sparse Fourier representation because of their spectral bias. This shows both these classes of functions can be compactly represented in the Fourier basis. The results here will become useful in the next section, where we present our main contribution on how we can leverage this compact representation, to precisely compute SHAP values cheaply.
3.1 Spectral bias of fully connected neural networks
The function a fully connected neural network represents at initialization is a sample from a Gaussian Process (GP) [Rasmussen, 2004] in the infinite-width limit. Here, the randomness is over the initial weights and biases. The kernel $K$ of this GP is called the Conjugate Kernel (CK) [Daniely et al., 2016, Lee et al., 2017]. As mentioned before, here we investigate the case where the inputs to the neural network are binary, i.e., $\mathcal{X}^{d}=\{0,1\}^{d}$, similar to [Yang and Salman, 2019, Valle-Perez et al., 2018]. The CK kernel Gram matrix formed on the whole input space $\mathcal{X}^{d}=\{0,1\}^{d}$, has a simple eigenvalue decomposition: $\mathcal{K}\in\mathbb{R}^{2^{d}\times 2^{d}},\mathcal{K}=\sum\limits_{f\in\{0,1\}^{n}}\lambda_{f}u_{f}u_{f}^{\top},$ where $u_{f}\in\mathbb{R}^{2^{d}}$ is the eigenvector formed by evaluating the Fourier basis function $\Psi_{f}$ for different values of $x\in\{0,1\}^{n}$. Moreover, Yang and Salman [2019] show a weak spectral bias result in terms of the degree of $f$. Namely, the eigenvalues corresponding to higher degree frequencies have smaller values  [^2] To be more precise, they show that the eigenvalues corresponding to even and odd degree frequencies form decreasing sequences. That is, even and odd degrees are considered separately.. Given the kernel, $K$, and viewing a randomly initialized neural network function evaluated on the boolean $\{0,1\}^{n}$ as a sample from a GP one can see the following: This sample, roughly speaking, looks like a linear combination of the eigenvectors with the largest eigenvalues. This is due to the fact that a sample of the GP can be obtained as $\sum\limits_{i=1}^{2^{n}}\lambda_{i}\bm{w_{i}}u_{i},\bm{w_{i}}\sim\mathcal{N}(0,1)$. Combining this with the spectral bias results implies that neural networks are low-degree functions when randomly initialized.
Going beyond neural networks at initialization, numerous studies have investigated the behavior of fully connected neural networks trained through (stochastic) gradient descent. Chizat et al. [2019], Jacot et al. [2018b], Du et al. [2018], Allen-Zhu et al. [2019a, b] found that the weights of infinite-width neural networks after training do not deviate significantly from their initialization, which has been dubbed "lazy training" by Chizat et al. [2019]. Lee et al. [2018, 2019] showed that training the last layer of an infinite-width randomly initialized neural network for an infinite amount of time corresponds to Gaussian process (GP) posterior inference with a certain kernel. Jacot et al. [2018b] extended the aforementioned results to training all the layers of a neural network (not just the final layer). They showed the evolution of an infinite-width neural network function can be described by the “Neural Tangent Kernel” [NTK, Jacot et al., 2018a] with the trained network yielding, on average, the posterior mean prediction of the corresponding GP after an infinite amount of training time. Lee et al. [2019] empirically showed the results carry over to the finite-width setting through extensive experiments. Yang and Salman [2019] again showed that the $u_{f}\in\mathbb{R}^{2^{d}}$ defined above are eigenvectors of the NTK Gram matrix and spectral bias holds. Gorji et al. [2023] validated these theoretical findings through extensive experiments in finite-width neural networks by showing that neural networks have “less tendency” to learn high-degree frequencies. We refer the reader to Appendix A.1 for a more comprehensive review.
The aforementioned literature shows that neural networks can be approximated by low-degree functions. By Proposition 2, they can be approximated by sparse functions for a large enough sparsity factor $k$. Our experiments in Section 5 provide further evidence that neural networks trained on real-world datasets are approximated well with sparse (and therefore compact) Fourier representations.
3.2 Sparsity of ensembles of decision trees
In our context, a decision tree is a rooted binary tree, where each non-leaf node corresponds to one of $n$ binary (zero-one) features, and each leaf node has a real number assigned to it. We denote the function a decision tree represents by $t:\{0,1\}^{n}\rightarrow\mathbb{R}$. Let $i\in[n]$ denote the feature corresponding to the root, and let $t_{\mathrm{left}}:\{0,1\}^{n-1}\rightarrow\mathbb{R}$ and $t_{\mathrm{right}}:\{0,1\}^{n-1}\rightarrow\mathbb{R}$ be the left and right sub-trees, respectively. Then the tree can be represented as:
$t(x)=\frac{1+(-1)^{\langle e_{i},x\rangle}}{2}t_{\mathrm{left}}(x)+\frac{1-(-1)^{\langle e_{i},x\rangle}}{2}t_{\mathrm{right}}(x)$ (2)
where, $e_{i}\in\{0,1\}^{n}$ is $i$’th indicator vector.
Thus, the Fourier transform of a decision tree can be computed recursively [Kushilevitz and Mansour, 1993, Mansour, 1994]. The degree of a decision tree function of depth $d$ is $d$, and if $|\text{supp}(t_{\mathrm{left}})|=k_{\mathrm{left}}$ and $|\text{supp}(t_{\mathrm{right}})|=k_{\mathrm{right}}$, then $|\text{supp}(t)|\leq 2(k_{\mathrm{left}}+k_{\mathrm{right}})$. As a result, a decision tree function is $k$-sparse, where $k=O(4^{d})$, although in some cases, when the decision tree is not balanced or cancellations occur, the Fourier transform can be sparser, i.e., admit a lower $k$, than the above upper bound on the sparsity suggests.
Due to the linearity of the Fourier transform, the Fourier transform of an ensemble of trees, such as those produced by the random forest, cat-boost [Dorogush et al., 2018], and XGBoost [Chen and Guestrin, 2016] algorithms/libraries, can be computed by taking the average of the Fourier transform of each tree. If the random forest model has $T$ trees, then its Fourier transform is $k=O(T4^{d})$-sparse and of degree $d$ equal to its maximum depth of the constituent trees.
4 Computing SHAP values with Fourier representation of functions
In the previous section, we saw that many real-world models trained on tabular/discrete data can be exactly represented (in the case of ensembles of decision trees), or efficiently approximated (in the case of neural networks) using a compact (sparse) Fourier representation. We saw neural networks have a tendency to learn approximately low-degree, and hence by Proposition 2, sparse functions. This has been attributed in numerous works to the reason why they generalize well and do not overfit despite their over-parameterized nature [Valle-Perez et al., 2018, Yang and Salman, 2019, Huh et al., 2022, Durvasula and Liter, 2020, Kalimeris et al., 2019, Neyshabur et al., 2017, Arpit et al., 2017]. We also saw that (ensembles) of decision trees, by nature, have sparse Fourier representations [Kushilevitz and Mansour, 1993, Mansour, 1994]. More generally, as made formal in Proposition 1 and the remarks after, any “simple” function that can be written as a summation of a “few” functions each depending on a “few” of the input variables is sparse and low-degree in the Fourier domain.
We propose the following method to approximate SHAP values for black-box functions. In the first step, given query access to a black-box function, we utilize a sparse Fourier approximation algorithm such as [Li and Ramchandran, 2015, Amrollahi et al., 2019, Li et al., 2015] to efficiently extract its sparse and hence compactly represented Fourier approximation. See Appendix A.2 for a more detailed explanation. Next, in our second step presented here, we use the Fourier representation to exactly compute SHAP values.
Let $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ be some predictor that we assume to be $k$-sparse. Assume $[n]\triangleq\{1,\dots,n\}$ as the set of features and let $x^{*}\in\{0,1\}^{n}$ be the instance we are explaining. Aligned with [Lundberg and Lee, 2017], we use $v_{h}(S)\triangleq\frac{1}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}h(x^{*}_{S},x_{[n]\setminus S})$ as the value function, which is the average prediction when fixing $x_{S}^{*}$ in the (background) dataset $\mathcal{D}$. We compute the Shapley values for a single Fourier basis function $\Psi_{f}(x)=(-1)^{\langle f,x\rangle},f\in\{0,1\}^{n}$ according to Equation 1:
$\phi_{i}^{\Psi_{f}}=\frac{1}{|\mathcal{D}|}\sum\limits_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}\cdot\sum\limits_{(x,y)\in\mathcal{D}}\left((-1)^{\langle f,x^{*}_{S\cup\{i\}}\oplus x_{[n]\setminus S\cup\{i\}}\rangle}-(-1)^{\langle f,x^{*}_{S}\oplus x_{[n]\setminus S}\rangle}\right)$ (3)
The $\oplus$ operator concatenates two vectors along the same axis.
This expression still has an exponential sum (in $n$), since we are summing over all subsets $S$. As a main contribution, we find a closed-form analytic expression for the inner summation using a combinatorial argument. This results in the following key Lemma:
Lemma 1.
Let $\Psi_{f}(x)=(-1)^{\langle f,x\rangle}$ be the Fourier basis function for some $f\in\{0,1\}^{n}$. Then the SHAP value of the Fourier basis function $\Psi_{f}$ with respect to the background dataset $\mathcal{D}$ is given as:
$\phi_{i}^{\Psi_{f}}=-\frac{2f_{i}}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}\mathmybb{1}_{x_{i}\neq x^{*}_{i}}(-1)^{\langle f,x\rangle}\frac{(|A|+1)\mod 2}{|A|+1}$ (4)
where $A\triangleq\{j\in[n]|x_{j}\neq x^{*}_{j},j\neq i,f_{j}=1\}$. (Proof in Appendix B.2)
Finally, by the linearity of SHAP values w.r.t. the explained function $h$, and by utilizing Lemma 1 we arrive at the final expression for SHAP values of $h$. We present this closed-form expression alongside its computational complexity in our main Theorem:
Theorem 2.
Let $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ be a $k$-sparse pseudo-boolean function with Fourier frequencies $f^{1},\dots,f^{k}\in\text{supp}(h)$ and amplitudes $\widehat{h}(f),\forall f\in\text{supp}(f)$. Let $\mathcal{D}$ be a background dataset of size $|\mathcal{D}|$. Then, Equation 5 provides a precise expression for the SHAP value vector $\phi^{h}=(\phi^{h}_{1},\dots,\phi^{h}_{n})$. One can compute this vector with $\Theta(n\cdot|\mathcal{D}|\cdot k)$ flops (floating point operations).
$\phi_{i}^{h}=-\frac{2}{|\mathcal{D}|}\sum\limits_{f\in\text{supp}(h)}\widehat{h}(f)\cdot f_{i}\cdot\sum\limits_{(x,y)\in\mathcal{D}}\mathmybb{1}_{x_{i}\neq x^{*}_{i}}(-1)^{\langle f,x\rangle}\frac{(|A|+1)\mod 2}{|A|+1}$ (5)
where $A$ is the same as in Lemma 1. (Proof in Appendix B.3)
Theorem 2 gives us a computationally efficient way to go from a Fourier representation of a function $h$ to SHAP values. The SHAP values from this equation are exact, i.e., as long as the Fourier representation is exact, the SHAP values are also precise values. This is in contrast to KernelSHAP, where the SHAP values are approximated using stochastic sampling and one needs to check for convergence to ensure the accuracy. The approximation in our method is constrained to computing the sparse Fourier representation, only in the black-box setting.
Most importantly, the sum in Equation 5 is tractable. This is because we overcome the exponential sum involved in Equation 1 by analytically computing the sum, with a combinatorial argument, for a single Fourier basis function. We show the number of flops that are required to compute SHAP values in Theorem 2 to be asymptotically equal to $\Theta(n\cdot|\mathcal{D}|\cdot k)$. We note that the $|\mathcal{D}|$ and $k$ factors in the asymptotic computational complexity arise from the two summations present in Equation 5. Through this expression, we are able to “linearize” the computation of SHAP values to a summation over the Fourier coefficients and background dataset. This allows us to maximally utilize the parallelization on multiple cores and/or GPUs to speed up the computation significantly. Therefore, in the presence of multiple cores or a GPU, we can get a speedup equal to the level of parallelization, as each core or worker can compute one part of this summation.
We implement our algorithm called FourierShap using JAX [Bradbury et al., 2018], which allows for fast vectorized operations on GPUs. Each term inside the summations of Equation 5 can be implemented with simple vector operations. Furthermore, summations over the $k$ different frequencies in the support of $h$ and also background data points can both be efficiently implemented and parallelized using this library using its vmap operator. We perform all upcoming experiments on a single GPU. Nevertheless, we believe faster computation can also be achieved by crafting dedicated code designed to efficiently compute Equation 5 on GPU.
Finally, we note that the function approximation (first step) is only done once, i.e., we compute the sparse Fourier approximation of the black-box only once. This is typically the most expensive part of the computation. For any new query to be explained, we resort to an efficient implementation of Equation 5. As our experiments will show, this yields orders of magnitudes faster computation than previous methods such as KernelShap where, as mentioned before in Section 2, each explained instance requires solving an expensive optimization problem.
5 Experiments
We assess the performance of our algorithm, FourierSHAP, on four different real-world datasets of varying nature and dimensionality. Three of our datasets are related to protein fitness landscapes [Poelwijk et al., 2019, Wu et al., 2016, Sarkisyan et al., 2016] and are referred to as “Entacmaea” (dimension $n=13$), “GB1” ($n=80$), and “avGFP” ($n=236$) respectively. The fourth dataset is a GPU-tuning [Nugteren and Codreanu, 2015] dataset referred to as “SGEMM” ($n=40$). The features of these datasets are binary (zero-one) and/or categorical with standard one-hot encodings. See Appendix C for dataset details.
For the Entacmaea and SGEMM datasets, we train fully connected neural networks with 3 hidden layers containing 300 neurons each. For GB1 we train ensembles of trees models of varying depths using the random forest algorithm and for avGFP we train again, ensembles of trees models of varying depths using the cat-boost algorithm/library [Dorogush et al., 2018].
Black-box setting. The first step of the FourierSHAP algorithm is to compute a sparse Fourier approximation of the black-box model. We use a GPU implementation of a sparse Walsh-Hadamard Transform (sparse WHT) a.k.a Fourier transform algorithm [Amrollahi et al., 2019] for each of the four trained models. The algorithm accepts a sparsity parameter $k$ which is the sparsity of the computed Fourier representation. Higher sparsity parameter $k$ results in a better function approximation but the sparse-WHT runtime increase linearly in $k$ as well. In Figure 1 we plot the accuracy of the Fourier function approximation as measured by the $R^{2}$-score for different values of $k$ (which result in different runtimes). The $R^{2}$ score is computed over a dataset formed by randomly sampling the Boolean cube $\{0,1\}^{n}$.
Figure 1: Step 1 of FourierShap: Accuracy of the Fourier transform (in approximating the black-box function) vs. runtime of the sparse Fourier algorithm. The accuracy is evaluated by $R^{2}$ score and comparing the outputs of the black-box and the Fourier representation on a uniformly generated random dataset on the Boolean cube $\{0,1\}^{n}$. For a fixed level of accuracy, higher depth trees require a higher number of Fourier coefficients $k$ therefore a higher runtime. For the case of trees, we eventually are able to reach a perfect approximation since the underlying function is truly sparse.
The second step of FourierSHAP utilizes the Fourier approximation to compute SHAP values using Equation 5. We implement this step using JAX library [Bradbury et al., 2018], and run it on a GPU. For each model to be explained, we choose four different values for the number of background samples and four different values for the number of query points to be explained, resulting in a total of 16 runs of FourierSHAP for each model. Error bars capture these variations. We take the runtime of KernelSHAP, with Github repo default settings, to be the base runtime all other methods are compared to, i.e., we assume its runtime is 1 unit.
Figure 2: Speedup vs. Accuracy. Speedup of different algorithms is reported as a multiple compared to the runtime of KernelShap. Accuracy is quantified by the $R^{2}$-score against ground truth SHAP values. DeepLift is a white-box algorithm for neural networks. LinRegShap is a black-box algorithm and a variance-reduced version of KernelShap. FastShap, a black-box algorithm, is a trained MLP to output SHAP values given inputs in one forward pass. We test MLPs of three different sizes for FastShap. FourierShap is ours. We are 10-10000x faster than LinRegShap on all dataset/model variations. More notably, we outperform DeepLift in the neural network model setting even though we assume only query access (black-box setting). We achieve higher accuracy than FastShap in 3/4 settings, while enabling a fine-grained control over the speed-accuracy trade-off.
In order to measure the accuracy of the SHAP values produced by our and other algorithms we need ground truth SHAP values. We use KernelSHAP to this extent. Note that KernelSHAP is inherently an approximation method for computing interventional SHAP values. However, this approximation becomes more precise by sampling more coalition subsets. Therefore, to generate ground truth values, for each dataset, we sample more and more coalition subsets and check for convergence in these values. A more detailed explanation can be found in Appendix E.3.
We compute the $R^{2}$ values of Shapley values computed by FourierSHAP (ours) vs. ground truth values to evaluate accuracy. For our method, a higher sparsity $k$ for the Fourier representation results in a more accurate function approximation therefore higher $R^{2}$ values for the SHAP value quality. On the other hand, a higher $k$ results in a slower runtime as Equation 5 is a sum over the $k$ different frequencies. In Figure 2 we plot this trade-off.
We compare against the following baselines in Figure 2. The first is LinRegShap, a variance-reduced version of KernelSHAP [Covert and Lee, 2020]. We found that although this algorithm requires fewer queries from the black-box, it takes orders of magnitudes longer to run compared to ours. Secondly, for the neural network models, we also compare against a state-of-the-art white-box method – DeepLift [Shrikumar et al., 2017]. This algorithm, requires access to the neural network’s activations in all layers. In comparison, we achieve a 10-100x speedup while being both more accurate and only assuming query access to the neural net (true black-box setting).
Finally, we compare against FastShap [Jethani et al., 2021], which is the closest to our setting, i.e., the model agnostic black-box, amortized setting. We provide higher accuracy for similar values of speedup compared to FastShap in 3/4 settings. More importantly though, we can see by controlling the sparsity parameter $k$ of the Fourier function approximation, we can control the tradeoff between accuracy and speed in a reliable and fine-grained manner. Whereas for FastShap the accuracy of the SHAP values relies on the functional approximation properties of the MLP which directly produces SHAP values. As seen in Figure 2, the MLP can behave in unpredictable ways. In this figure we can see that increasing the depth of the model (and hence the approximation capability of the MLP) does not have a reliable effect on SHAP accuracy. Finally, in FastShap for a different choice of background dataset a new MLP model has to be trained from scratch. Since, the MLP approximates the process of directly computing SHAP values. Whereas we can support different sets of background datasets since our Fourier functional approximation is done on the black box predictor, and not the whole process of producing SHAP values in an end-to-end manner (FastShap).
Tree setting. FourierSHAP can also be utilized for the computation of SHAP values for (ensembles of) trees in the white-box setting, i.e., where full access to the tree’s structure is available. In this setting, the exact sparse Fourier representation of an ensemble of trees can be efficiently computed (the first step of FourierSHAP) using Equation 2. With the exact Fourier representation at hand, SHAP values can be efficiently and exactly computed using Equation 5, the second step of FourierSHAP. This makes our method a highly parallelizable alternative for TreeShap [Lundberg et al., 2020] with the potential for order of magnitude speedups.
To demonstrate our method’s ability in fast computation of SHAP values in this setting, we compute SHAP values for random forests fitted on all aforementioned real-world datasets. We compare FourierSHAP’s speedup over TreeSHAP [Lundberg et al., 2020], to FastTreeSHAP [Yang, 2021], a fast implementation of TreeSHAP, the GPU implementation of TreeSHAP [Mitchell et al., 2022b] and PLTreeSHAP [Zern et al., 2023]. To the best of our knowledge, these are the fastest available frameworks for computation of the “interventional” SHAP values. Figure 3 shows the superior speed of our method over all state-of-the-art algorithms in most settings, which weakens by the increment of depth and number of features, resulting in more frequencies and computation overhead. Note that the SHAP values computed by all methods are precise and identical to the values produced by TreeSHAP, which is to be expected for FourierSHAP, given access to exact Fourier representation of random forests.
Figure 3: Speedup vs. depth of tree, for different algorithms, reported as a multiple compared to the runtime of TreeSHAP. FourierSHAP is ours. As other baselines we have a GPU implementation of TreeSHAP, FastTreeSHAP, and PLTreeSHAP. We achieve order of magnitude speedups over all depths on the Entacmaea and SGEMM datasets. We also achieve significant speedups in the other two datasets; however, the edge diminishes as the maximum depth increases.
Conclusions
We illustrated in theory and practice how many black-box functions can be represented or efficiently approximated by a compact Fourier representation. We proved that SHAP values of Fourier basis functions admit a closed form expression, and therefore, we can compute SHAP values efficiently using the compact Fourier representation. Moreover, this closed form expression is amenable to parallelization. These two factors helped us in gaining speedups of 10-10000x over baseline methods for the computation of SHAP values.
Appendix A Relevant work
A.1 Neural network theory and simplicity biases
With regards to simplicity biases in fully connected neural networks, a substantial amount of research has been dedicated to analyzing neural networks in function space. This line of research is dedicated to firstly showing that “infinite-width”, randomly initialized (with Gaussian distribution) neural networks are distributed as Gaussian Processes (GPs) and secondly computing the kernel associated to the GP [Neal, 1996, Williams, 1996, Cho and Saul, 2009, Hazan and Jaakkola, 2015, Lee et al., 2017, Ancona et al., 2018, Daniely, 2017]. The kernel associated with the GP is commonly known as the “conjugate kernel” [Daniely, 2017] or the “NN-GP kernel”[Lee et al., 2017]. Other works show that in infinite-width neural networks weights after training via SGD do not end up too far from the original [Chizat et al., 2019, Jacot et al., 2018a, Du et al., 2018, Allen-Zhu et al., 2019a, b], referred to as “lazy training” by Chizat et al. [2019]. This allowed Jacot et al. [2018a] to prove that the evolution of an infinite-width neural network during training can be described by a kernel called the “Neural Tangent Kernel”. Lee et al. [2019] showed, through extensive experiments that the same behavior holds even for the more realistic case of neural nets of finite width. Empirically speaking, it was shown by Rahaman et al. [2019] that a neural net with one input tends to learn sinusoids of lower frequencies in earlier epochs than those with higher frequencies. By analyzing the spectrum of the NTK’s Gram matrix, Ronen et al. [2019], Basri et al. [2020] were able to formally prove this empirical finding. Yang and Salman [2019], Fan and Wang [2020] analyze the spectra of the NTK gram matrix for higher dimensional inputs. Specifically, Yang and Salman [2019], Valle-Perez et al. [2018] provide simplicity bias results for the case where the inputs to the neural net are Boolean (zero-one) vectors.
A.2 Sparse and low-degree Fourier transform algorithms
We now discuss algorithms that efficiently approximate general black-box predictors by a Fourier sparse representation. Let $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ be a any function. We assume we have query access to $h$. That is, we can arbitrarily pick $x\in\{0,1\}^{n}$ and query $h$ for its value $h(x)$. Without any further assumptions, computing the Fourier transform requires us to query exponentially, to be precise $2^{n}$, many queries: one for every $x\in\{0,1\}^{n}$. Furthermore, classical Fast Fourier Transform (FTT) algorithms are known to take at least $\Omega(2^{n}\log(2^{n}))$ time.
Under the additional assumption that $h$ is $k$-sparse, works such as Cheraghchi and Indyk [2017], Amrollahi et al. [2019], Scheibler et al. [2015], Kushilevitz and Mansour [1993], Li and Ramchandran [2015], Li et al. [2015] provide algorithms that obtain the Fourier transform more efficiently. In particular, Amrollahi et al. [2019] provide algorithms with query complexity $O(nk)$ and time complexity $O(nk\log k)$ time. Assuming further that the function is of degree $d=o(n)$, the query complexity reduces to $O(kd\log n)$, with run time still polynomial in $n,k,d$. Crucially, even if the function $h$ is not $k$-sparse, Algorithm RobustSWHT of Amrollahi et al. [2019] yields the best $k$-sparse approximation in the $\ell_{2}-\ell_{2}$ sense. More precisely, let us denote by $h_{k}:\{0,1\}^{n}\rightarrow$ the function that is formed by only keeping the top $k$ non-zero Fourier coefficients of $h$ and setting the rest to zero. Then the algorithm returns a $O(k)$-sparse function $g$ such that:
$\sum\limits_{f\in\{0,1\}^{n}}(\widehat{g}(f)-\widehat{h}(f))^{2}\leq C(1+\epsilon)\min\limits_{\text{all }k\text{-sparse }g}\sum\limits_{f\in\{0,1\}^{n}}(\widehat{g}(f)-\widehat{h_{k}}(f))^{2},$
where $C$ is a universal constant. By Parseval’s identity, the same holds if the summations were over the time (input) domain instead of the frequency domain.
A.3 Shapley values in the context of Machine learning
In the context of ML, many works have derived a different notion of Shapley value depending on what they mean by data distribution, deleted features, etc. We refer the reader to the survey by Sundararajan and Najmi [2020], Janzing et al. [2020] for a comprehensive overview. In this work we focus on the notion of SHAP introduced by Lundberg and Lee [2017], Lundberg et al. [2020] also known as “Interventional” [Janzing et al., 2020, Van den Broeck et al., 2021] or “Baseline” SHAP [Sundararajan and Najmi, 2020] where the missing features are integrated out from the marginal distribution as opposed to the conditional distribution (see Section 2). As pointed out by Janzing et al. [2020] there are two main ways to define the SHAP value “interventional” and “observational” SHAP. These are referred to by Sundararajan and Najmi [2020] as “baseline” and “conditional” SHAP respectively.
As pointed out by Janzing et al. [2020] the difference between these definitions can be better viewed with the lens of causality [Pearl, 2009]. Roughly speaking “observational” SHAP tells us about how influential a feature is to the prediction of the predictor if it goes from the state of being unobserved to observed. “Interventional” SHAP is causal and tells us how influential a feature is if we were to reach in (through a process called an intervention) and change that feature in order to change the prediction. Put into the context of credit scores and loan approvals, “observational” SHAP will provide us with important features which are “observed” by the predictor and hence are influential in predicting if a particular loan request will be rejected or approved. Interventional SHAP would tell us which feature we could change or “intervene” in order to change the outcome of the loan request.
The original (ML) SHAP paper [Lundberg and Lee, 2017] proposes “observational” SHAP as the correct notion of SHAP. Van den Broeck et al. [2021], Arenas et al. [2021] provide intractability results for observational SHAP in a variety of simple distributional assumptions on the data and simple predictors $f$. This has lead to many attempts to approximate observational SHAP values [Lundberg et al., 2020, Covert and Lee, 2020, Aas et al., 2021, Kwon et al., 2021, Sundararajan and Najmi, 2020]. It is interesting to note that the version of “Kernel”-SHAP in Lundberg and Lee [2017] is also an approximation for observational SHAP values that ends up coinciding precisely with interventional SHAP values, which explains a lot of the confusion in the community. Janzing et al. [2020] boldly claims that researchers should stop the pursuit of approximations to “observation” SHAP values as it lacks certain properties, for example, sensitivity i.e. the SHAP value of a feature can be non-zero while the predictor $f$ has no dependence on that feature. This phenomenon happens because when features are correlated, the presence of a feature can provide information about other features that the predictor does depend on. This does not happen in interventional SHAP. Finally, Chen et al. [2020] argues that both SHAP definitions are worthy of pursuit. They emphasize that the interventional framework provides explanations that are more “true to the model”, and the observational approach’s explanations are more “true to the data”.
Appendix B Proofs
Before we start with the proofs we review the Fourier analysis and synthesis equations. As we mentioned in the Background Section 2, the Fourier representation of the pseudo-boolean function $h:\{0,1\}^{n}\rightarrow\mathbb{R}$ is the unique expansion of $h$ as follows:
$h(x)=\frac{1}{\sqrt{2^{n}}}\sum\limits_{f\in\{0,1\}^{n}}\widehat{h}(f)(-1)^{\langle f,x\rangle}$
This is the so called Fourier “synthesis” equation.
The Fourier coefficients $\widehat{h}(f)$ are computed by the Fourier “analysis” equation:
$\widehat{h}(f)=\frac{1}{\sqrt{2^{n}}}\sum\limits_{x\in\{0,1\}^{n}}h(x)(-1)^{\langle f,x\rangle}$ (6)
B.1 Proof of propositions
We can now prove Proposition 1: See 1
Proof.
Let $g:\{0,1\}^{n}\rightarrow\mathbb{R}$ be a function dependent on exactly $d$ variables $x_{i_{1}},\dots,x_{i_{d}}$, where $i_{1},\dots,i_{d}\in[n]$ are distinct indices. We show that for any frequency $f\in\{0,1\}^{n}$, if $f_{j}=1$ for some $j\notin S\triangleq\{i_{1},\dots,i_{d}\}$, then, $\widehat{g}(f)=0$. From the Fourier analysis Equation 6 we have:
$\widehat{g}(f)$ $=\frac{1}{\sqrt{2^{n}}}\sum\limits_{x\in\{0,1\}^{n}}g(x)(-1)^{\langle f,x\rangle}=\sum\limits_{x_{i}:i\in S}\sum\limits_{x_{j}:j\in[n]\setminus S}g(x)(-1)^{\langle f_{S},x_{S}\rangle}(-1)^{\langle f_{[n]\setminus S},x_{[n]\setminus S}\rangle}$
=$\stackrel{{\scriptstyle\textnormal{(i)}}}{{\mathstrut{=}}}\sum\limits_{x_{i}:i\in S}g(x)(-1)^{\langle f_{S},x_{S}\rangle}\sum\limits_{x_{j}:j\in[n]\setminus S}(-1)^{\langle f_{[n]\setminus S},x_{[n]\setminus S}\rangle}\stackrel{{\scriptstyle\textnormal{(ii)}}}{{\mathstrut{=}}}\sum\limits_{x_{i}:i\in S}g(x)(-1)^{\langle f_{S},x_{S}\rangle}\cdot 0=0$
Where Equation B.1 holds because $g$ is only dependent on the variables in $S$ and Equation B.1 holds by checking the inner sum has an equal number of $1$ and $-1$ added together.
The proof of the proposition follows by the linearity of the Fourier transform. ∎
Moving on to Proposition 2: See 2
Proof.
We simply note that the number of frequencies $f\in\{0,1\}^{n}$ of degree at most $d$ is equal to $\sum\limits_{i=0}^{d}\binom{n}{i}$. This sum is $O(n^{d})$ for $d$ constant w.r.t $n$. ∎
B.2 Proof of Lemma 1
Proof.
We start from Equation 3:
$\phi_{i}^{\Psi_{f}}$ $=\sum\limits_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}\cdot\frac{1}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}\left((-1)^{\langle f,x^{*}_{S\cup\{i\}}\oplus x_{[n]\setminus S\cup\{i\}}\rangle}-(-1)^{\langle f,x^{*}_{S}\oplus x_{[n]\setminus S}\rangle}\right)$
$=\frac{1}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}\sum\limits_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}(-1)^{\langle f_{-i},x^{*}_{S}\oplus x_{[n]\setminus S\cup\{i\}}\rangle}\left((-1)^{f_{i}x^{*}_{i}}-(-1)^{f_{i}x_{i}}\right)$
By checking all 8 possible combinations of $x_{i},x_{i}^{*},f_{i}\in\{0,1\}$, one can see that $(-1)^{f_{i}x^{*}_{i}}-(-1)^{f_{i}x_{i}}=2f_{i}(x_{i}-x_{i}^{*})$. This is simply because the two exponents only differ when $f_{i}=1$ and $x_{i}\neq x_{i}^{*}$.
To determine $(-1)^{\langle f_{-i},x^{*}_{S}\oplus x_{[n]\setminus S\cup\{i\}}\rangle}$, we partition $[n]\setminus\{i\}$ into two subsets $A\triangleq\{j\in[n]|x_{j}\neq x^{*}_{j},j\neq i,f_{j}=1\}$ and $B\triangleq[n]\setminus A\cup\{i\}$. Doing this, we can factor out $(-1)^{\langle f_{-i},x_{-i}\rangle}$ and determine the rest of the sign based on the number of indices in $S$ where $x$ and $x^{*}$ disagree and $f_{i}=1$. This is equal to $|A\cap S|$:
$\phi_{i}^{\Psi_{f}}=\frac{2f_{i}}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}(x_{i}-x^{*}_{i})(-1)^{\langle f_{-i},x_{-i}\rangle}\sum\limits_{S\subseteq[n]\setminus\{i\}}\frac{|S|!(n-|S|-1)!}{n!}(-1)^{|A\cap S|}$
$A$ and $B$ partition $[n]\setminus\{i\}$, therefore we split the inner sum as follows:
$\phi_{i}^{\Psi_{f}}=\frac{2f_{i}}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}(x_{i}-x^{*}_{i})(-1)^{\langle f_{-i},x_{-i}\rangle}\sum\limits_{\tilde{B}\subseteq B}\sum\limits_{\tilde{A}\subseteq A}\frac{(|\tilde{A}|+|\tilde{B}|)!(n-(|\tilde{A}|+|\tilde{B}|)-1)!}{n!}(-1)^{|\tilde{A}|}\$
Since the inner expression only depends on the cardinalities of $\tilde{A}$ and $\tilde{B}$ we can recast the inner sum to be over numbers instead of subsets by counting the number of times each cardinality appears in the summation:
$\phi_{i}^{\Psi_{f}}$ $=\frac{2f_{i}}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}(x_{i}-x^{*}_{i})(-1)^{\langle f_{-i},x_{-i}\rangle}\sum_{b=0}^{n-|A|-1}\sum_{a=0}^{|A|}\binom{n-|A|-1}{b}\binom{|A|}{a}\frac{(a+b)!(n-a-b-1)!}{n!}(-1)^{a}$
$=\frac{2f_{i}}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}(x_{i}-x^{*}_{i})(-1)^{\langle f_{-i},x_{-i}\rangle}\sum_{a=0}^{|A|}(-1)^{a}\sum_{b=0}^{n-|A|-1}\frac{\binom{|A|}{a}\binom{n-|A|-1}{b}}{n\binom{n-1}{a+b}}$ (7)
Now we find a closed-form expression for the innermost summation in the above Equation, which is a summation over $b$ where $a$ is fixed:
$\sum_{b=0}^{n-|A|-1}\frac{\binom{|A|}{a}\binom{n-|A|-1}{b}}{n\binom{n-1}{a+b}}$ =$\stackrel{{\scriptstyle\textnormal{(iii)}}}{{\mathstrut{=}}}\sum_{b=0}^{n-|A|-1}\frac{\binom{|A|}{a}\binom{n-|A|-1}{b}}{n\frac{\binom{n-1}{|A|}}{\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}}\binom{|A|}{a}\binom{n-|A|-1}{b}}$
$=\frac{1}{n\binom{n-1}{|A|}}\sum_{b=0}^{n-|A|-1}\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}$
=$\stackrel{{\scriptstyle\textnormal{(iv)}}}{{\mathstrut{=}}}\frac{1}{n\binom{n-1}{|A|}}\binom{n}{|A|+1}$
$=\frac{1}{|A|+1}$ (8)
In Equation B.2, we use the following identity: $\binom{n-1}{a+b}\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}=\binom{n-1}{|A|}\binom{|A|}{a}\binom{n-|A|-1}{b}$. This can be checked algebraically by simply writing down each binomial term as factorials and doing the cancellations:
$\binom{n-1}{a+b}\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}$
$=\frac{(n-1)!}{(a+b)!(n-a-b-1)!}\cdot\frac{(a+b)!}{a!b!}\cdot\frac{(n-a-b-1)!}{(|A|-a)!(n-a-b-1-(|A|-a))!}$
$=\frac{(n-1)!}{a!b!(|A|-a)!(n-|A|-b-1)!}.$
$=\frac{(n-1)!}{|A|!(n-|A|-1)!}\cdot\frac{|A|!}{a!(|A|-a)!}\cdot\frac{(n-|A|-1)!}{b!(n-|A|-1-b)!}$
$=\binom{n-1}{|A|}\binom{|A|}{a}\binom{n-|A|-1}{b}$
In Equation B.2, we use $\sum_{b=0}^{n-|A|-1}\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}=\binom{n}{|A|+1}$ which holds because of the following double-counting argument. The term $\binom{n}{|A|+1}$ counts the number of ways to choose a subset of size $|A|+1$ from a set of $n$ elements. Imagine elements are numbered from 1 to n, and $\pi_{1}<\pi_{2}<...<\pi_{|A|+1}\in[n]$ represent $|A|+1$ chosen elements. Let’s condition on $\pi_{a+1}=(a+b+1)$ where possible values for $b$ can only be $10\leq b\leq n-|A|-$ 1. This is due to the fact that $\pi_{a+1}<a+1$ implies $\pi_{1},...,\pi_{a}$ are chosen from less than $a$ elements, and similarly $\pi_{a+1}>n-|A|+a$ implies $\pi_{a+2},...,\pi_{|A|+1}$ ($|A|-a$ chosen elements) are chosen from less than $|A|-a$ elements, which are both impossible. Given the condition, $a$ elements numbered lower than $(a+b+1)$ and $|A|-a$ elements numbered larger than $(a+b+1)$ are also chosen. This is possible in $\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}$ ways. Therefore, keeping in mind that $a$ is fixed, $\sum_{b=0}^{n-|A|-1}\binom{a+b}{a}\binom{n-a-b-1}{|A|-a}$ should give the number of ways to choose a subset of size $|A|+1$ from a set of $n$ elements, through the new perspective.
Based on Equation 8, we see that the innermost summation in Equation B.2 is only dependent on $|A|$. Thus, we rewrite Equation B.2 as follows:
$\phi_{i}^{\Psi_{f}}$ $=\frac{2f_{i}}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}(x_{i}-x^{*}_{i})(-1)^{\langle f_{-i},x_{-i}\rangle}\frac{(|A|+1)\mod 2}{|A|+1}$
By absorbing the sign of $(x_{i}-x_{i}^{*})$ into the $(-1)^{{\langle f_{-i},x_{-i}\rangle}}$ term we arrive at Equation 4:
$\phi_{i}^{\Psi_{f}}=-\frac{2f_{i}}{|\mathcal{D}|}\sum\limits_{(x,y)\in\mathcal{D}}\mathmybb{1}_{x_{i}\neq x^{*}_{i}}(-1)^{\langle f,x\rangle}\frac{(|A|+1)\mod 2}{|A|+1}$
B.3 Proof of Theorem 2
Proof.
Proof of Equation 5 simply follows from the fact that SHAP values are linear w.r.t. the explained function. Regarding the computational complexity we restate Equation 5
$\phi_{i}^{h}=-\frac{2}{|\mathcal{D}|}\sum\limits_{f\in\text{supp}(h)}\widehat{h}(f)\cdot f_{i}\sum\limits_{(x,y)\in\mathcal{D}}\mathmybb{1}_{x_{i}\neq x^{*}_{i}}(-1)^{\langle f,x\rangle}\frac{(|A|+1)\mod 2}{|A|+1}$
where $A\triangleq\{j\in[n]|x_{j}\neq x^{*}_{j},j\neq i,f_{j}=1\}$.
We first do a pre-processing step for amortizing the cost of computing $|A|$: we compute $\tilde{A}\triangleq\{j\in[n]|x_{j}\neq x^{*}_{j},f_{j}=1\}$ which takes $\Theta(|\mathcal{D}|n)$ flops.
We assume we are computing the whole vector $\Phi^{h}=(\Phi^{h}_{1},\dots,\Phi^{h}_{n})$, that is we are compute SHAP values for all $i\in[n]$ at the same time. Going back to the inner summation above, computing $A$ (and $|A|$) for different values of $i\in[n]$ is $\Theta(n)$ if we utilize the pre-computed $\tilde{A}$. The inner product $\langle f,x\rangle$ is not dependent on $i$ and is $\Theta(n)$ flops. Computing $\mathmybb{1}_{x_{i}\neq x^{*}_{i}}$ for different values of $i\in[n]$ is $\Theta(n)$. Therefore, the inner expression of the summand takes $\Theta(n)$ flops for a fixed data-point $x\in\mathcal{D}$ and $f\in\text{supp}(h)$.
Computing the inner sum for any fixed frequency $f\in\text{supp}(h)$ is $\Theta(n|\mathcal{D}|)$, because we are summing over $|\mathcal{D}|$ vectors each of size $n$ (the vector which holds SHAP value for each $i\in[n]$). Moving on to the outer sum each evaluation of the inner sum is $\Theta(n|\mathcal{D}|)$ and it results in a vector of size $n$ (one element for each SHAP value). The multiplication of $\widehat{h}(f)\cdot f_{i}$ is $\Theta(n)$. Therefore the cost of the inner sum dominates i.e. $\Theta(n|\mathcal{D}|)$. Since we are summing over the whole support the total number of flops is: $\Theta(n|\mathcal{D}|k)$ where $k=|\text{supp}(h)|$. ∎
Appendix C Datasets
We list all the datasets used in the Experiments Section 5.
Entacmaea quadricolor fluorescent protein. (Entacmaea)
Poelwijk et al. [2019] study the fluorescence brightness of all $2^{13}$ distinct variants of the Entacmaea quadricolor fluorescent protein, mutated at $13$ different sites.
GPU kernel performance (SGEMM).
Nugteren and Codreanu [2015] measures the running time of a matrix product using a parameterizable SGEMM GPU kernel, configured with different parameter combinations. The input has $14$ categorical features. After one-hot encoding the dataset is $40$-dimensional.
Immunoglobulin-binding domain of protein G (GB1).
Wu et al. [2016] study the “fitness” of variants of protein GB1, that are mutated at four different sites. Fitness, in this work, is a quantitative measure of the stability and functionality of a protein variant. Given the $20$ possible amino acids at each site, they report the fitness for $00020^{4}=160,$ 000 possible variants, which we represent with one-hot encoded 80-dimensional binary vectors.
Green fluorescent protein from Aequorea victoria (avGFP).
Sarkisyan et al. [2016] estimate the fluorescence brightness of random mutations over the green fluorescent protein sequence of Aequorea victoria (avGFP) at 236 amino acid sites. We transform the amino acid features into binary features indicating the absence or presence of a mutation at each amino acid site. This converts the original $54,024$ distinct amino acid sequences of length 236 into $49,089$ 236-dimensional binary data points.
Appendix D FourierShap Implementation
We implemented four versions of our algorithm, i.e., Equation 5, using Google JAX library [Bradbury et al., 2018]. JAX provides a flexible framework for developing high-performace functions for vectorized computations. JAX enables automatic performance optimisation of algebraic computation as well as just-in-time (JIT) compilation of the vectorized functions for faster iterations at runtime.
The full code base is provided as supplementary material to the paper. We refer the reader there for more details. Here we give a high level overview of the four versions we implemented and experimented with:
  • Base: In this version, frequencies are represented as $n$-dimensonal binary vectors. A simple implenetation is provided for computing the SHAP values given a single frequency, a single background instance, and a single query instance. This is next extended to multiple frequencies, multiple background instances, and multiple query instances using multiple JAX vmaps.
  • Precompute: This version is a modified version of the Base version. We take a closer look at Equation 5 and pre-compute terms dependent only on frequencies $f$ and background dataset points $x$ (not including terms dependent on the query point $x^{*}$). These are then loaded from memory at run-time. This version was faster than Base in all of our experiments, but inherently requires more memory; in our case, more GPU memory.
  • Sparse: Here we utilize the sparsity of frequencies $f\in\{0,1\}^{n}$ i.e. the fact that frequencies have mostly zero entries in practice. In this version, for each frequency $f$, we focus on positions $i$ where $f_{i}=1$. Inspecting Equation 5, these are the only positions where $f$ can affect the final SHAP value vector.
  • Positional: This version brings again builds on the previous and computes SHAP values separately for each coordinate $i\in[n]$. This adds one extra precomputation step to map frequencies to coordinates they affect, but enables mathematical simplifications as well as potential for extra vectorization. We observed that although this extra precomputation step could become time-consuming for large set of frequencies, it can result in a significantly faster SHAP value computation at run-time in the case of low-degree Fourier spectrums, compared to the Sparse version.
Appendix E Experiment Details
The code for running the experiments and the implementations of all modules will be open-sourced once the double-blind review process is over. We run all experiments on a machine with one NVIDIA GeForce RTX 4090 GPU, on servers with Intel(R) Xeon(R) CPU E3-1284L v4 @ 2.90GHz, restricting the memory/RAM to 20 GB, which was managed with Slurm.
E.1 Black-box
For the Entacmaea and SGEMM datasets, we train fully connected neural networks with 3 hidden layers containing 300 neurons each. The network is trained using the means-squared loss and ADAM optimizer with a learning rate of 0.01. For GB1 we train ensembles of trees models of varying depths using the random forest algorithm using the sklearn library [Pedregosa et al., 2011]. For avGFP we train again, ensembles of trees models with 10 trees of varying depths using the cat-boost algorithm/library[Dorogush et al., 2018]. All other setting are set to the default in both cases. Model accuracy for different depths are plotted in Figure 4.
Figure 4: Model accuracy of tree models evaluated on the test set for different depths
For each dataset, we use all possible combination of number of background samples and query instances (instances to be explained) from $\{10,20,30,40\}$, resulting in a total of 16 runs per dataset. Error bars in Figure 2 capture the variation in speedups.
E.1.1 FourierSHAP
For the first step of the FourierSHAP, i.e., computing a sparse Fourier approximation of the black-box model, we use an implementation of a sparse Walsh-Hadamard Transform (sparse WHT) a.k.a Fourier transform algorithm [Amrollahi et al., 2019] for each of the four trained models. The implementation is done using Google JAX [Bradbury et al., 2018] library as part of this work, and is available in the project repository.
For the second step of FourierSHAP, i.e., utilizing the approximated Fourier spectrum to compute SHAP values as per Equation 5, we use our “Precompute” implementation. We discussed the details in the previous Section, Section D.
To showcase the flexibility of our method in controlling the trade-off between speed and accuracy in Figure 2, we prune the computed sparse Fourier spectrum in the first step, and remove frequencies with amplitudes smaller than a specific threshold from the spectrum. For the threshold, we use $10$ values between $0.0001$ and $0.05$ and specified a minimal descriptive subset of them as points in Figure  2.
E.1.2 KernelShap
For KernelShap, we use the standard library provided by the writers of the paper [Lundberg and Lee, 2017]  [^3] https://github.com/slundberg/shap with its default settings. As part of the default setting, the paired sampling trick is also enabled, which shown to be beneficial for faster convergence of KernelSHAP. KernelShap is written in C and is to our knowledge the fastest implementation of this algorithm.
To ensure a fair comparison, we run models on GPU wherever possible, i.e., for Neural Networks on Entacmaea and SGEMM datasets.
E.1.3 Fast-SHAP
We tried our best to capture the full potential of FastSHAP in predicting SHAP values in terms of speed and accuracy, to ensure a practical and grounded comparison to our method. Here are the details on training Fast-SHAP as a baseline for computing interventional SHAP values:
  • Imputer: "Imputer" is a FastSHAP module used in the computation of neural networks’ loss, acting as the value function in SHAP formula, that generates model’s prediction using a strategy for treating features excluded in the subset, given the predictor and a subset of features. To compute interventional SHAP values, we use MarginalImputer, implemented in the original FastSHAP repo, which computes mean predictions when using the backgorund dataset’s values for excluded features. Therefore, each trained FastSHAP model is specific to a fixed background dataset, as the Imputer used in its loss is. In our experiments, we use four different background datasets with multiple sizes per (real-world) dataset, which lead to training four FastSHAP models per setting.
  • Feature Subset Sampling: We train FastSHAP models with $1$, $4$, and $16$ feature subset samples per input. Although we did not observe monotonic improvements by increasing the number of feature subset samples, we decided to use the models trained with $16$ samples per input to compare our method with, as it was mostly performing the best in terms of accuracy.
  • Paired Sampling: We enable paired sampling in FastSHAP, which is a trick to pair each feature subset sample $s\in\{0,1\}^{n}$ with its complement $1-s$, that is shown to be beneficial in reducing the variance and improving the accuracy, in both KernelSHAP and FastSHAP.
  • Neural Network Architectures: We use MLPs with three different sizes to train FastSHAP, and reported the results for each separately:
    • Small (2-layer): $in\times 128\times out$.
    • Medium (4-layer): $in\times 128\times 128\times 128\times out$.
    • Large (6-layer): $in\times 128\times 128\times 128\times 128\times 128\times out$.
    We use ReLU as the activation function in all models.
  • Hyperparameters: We use the training batch size of $64$ and FastSHAP’s defaults for all other components and hyper-parameters. We train each model with early stopping and up to $200$ epochs.
Table 1 shows the time required to train the FastSHAP models used in our black-box experiments. Both FourierShap (ours) and FastSHAP are amortized methods that have a heavier “pre-computation” step. For FastShap this pre-computation appears as training an MLP that directly predicts SHAP values and for us this appears as computing a Fourier transform. When comparing FastShap’s initial training time to ours (reported in Figure 1), we can see our method has a considerably lower pre-computation time.
Dataset Black-box model FastSHAP size Training time (m) for background dataset size Total training time (m)
10 20 30 40
Entacmaea MLP Small 5 8 15 18 46
Medium 2 4 6 10 22
Large 3 6 11 11 31
SGEMM MLP Small 47 130 249 173 599
Medium 55 52 162 128 397
Large 72 73 152 91 388
GB1 Random Forest (depth=5) Small 114 148 42 173 477
Medium 83 207 55 130 475
Large 54 102 155 184 495
avGFP Catboost (depth=5) Small 24 28 36 39 127
Medium 35 58 41 62 196
Large 11 20 22 25 78
Table 1: Training times of FastSHAP models used in our black-box experiments (in minutes). Unlike FourierSHAP (ours), FastSHAP needs to be separately trained for each background dataset. Part of the difference in the training times are due to the variance in the number of training epochs before the early stopping occurs. Original datasets are also of varying sizes that lead to different number of training samples per epoch.
E.1.4 LinRegShap
LinRegShap, is a variance-reduced version of KernelSHAP [Covert and Lee, 2020]. We again use the implementation of the original writers [^4] https://github.com/iancovert/shapley-regression. Their implementation includes automatic detection of the convergence of stochastic sampling which is meant to speed up the algorithm by taking less samples from the black box. Furthermore, as per the default setting, we also allowed paired sampling.
To ensure a fair comparison, we run models on GPU wherever possible, i.e., for Neural Networks on Entacmaea and SGEMM datasets.
E.1.5 DeepLift
For DeepLift we as well use the library of Lundberg and Lee [2017]  [^5] https://github.com/slundberg/shap with default settings. In this setting the neural network is passed to the algorithm on a GPU to make sure this algorithm is as fast as possible.
E.2 Trees
We fit random forests of maximum depths ranging from 3 to 8 with $20$ estimators on $90\%$ of all four datasets used in the black-box setting. We compare performance of four algorithms in computing SHAP values for these random forest models; the classic TreeSHAP as well as its GPU implementation  [^6] https://github.com/slundberg/shap, FastTreeSHAP  [^7] https://github.com/linkedin/FastTreeSHAP, and our FourierSHAP. We always use $100$ datapoints as the background data samples, and $100$ datapoints to explain and compute the SHAP values for. We report the speedup of each algorithm over classic TreeSHAP in Figure 3, with error bars showing the standard deviation in speedup over five independent runs.
For the first step of the FourierSHAP, we derive the exact sparse Fourier representation from the ensemble of trees, accessing the tree structures and using Equation 2. We also perform a pruning on the exact sparse Fourier representation derived from the ensemble of trees, and keep the frequencies with largest amplitudes that cover at least $99.95\%$ of the original Fourier spectrum’s energy. We also make sure to only remove frequencies with apmlitudes smaller than $0.005$.
To run FourierSHAP, we use our “Precompute” implementation for Entacmaea, and “Positional” version for other datasets, given larger feature spaces and the low-degreeness of the spectrum as a result of bounded maximum depth. See Section D for implementation details. We compare the resulted SHAP values with TreeSHAP results and achieve $R^{2}$ of at least $0.99$, ensuring precise values computed by FourierSHAP.
E.3 Using KernelShap to produce ground truth SHAP values
As mentioned before, in order to measure the accuracy of the SHAP values produced by our and other algorithms we need ground truth SHAP values. We use KernelShap to this extent. KernelShap is an approximation method for computing interventional SHAP values. This approximation becomes more precise by sampling more coalition subsets. Therefore, to generate ground truth values, for each dataset, we sample more and more and check for convergence in these values.
Figure 5: From left to right, 1- SHAP values generated by KernelSHAP, 2- SHAP values generated by FourierSHAP (ours), 3- and the ground truth SHAP values computed by the original exponential SHAP formula Equation1 on the Entacmaea dataset. In 3, Computation of ground truth SHAP values using the exponential formula is possible due to the dataset containing all $2^{13}$ possible boolean feature vectors. 10 query points and a background dataset points are chosen at random and are of size 10. This figure shows that both KernelSHAP and our method compute exact ground truth SHAP values.
In order to check if KernelShap values have converged to ground truth interventional SHAP values, we update the number of subset samples KernelShap uses to generate SHAP values for each instance. Looking into KernelShap repo  [^8] https://github.com/shap/shap/blob/master/shap/explainers/_kernel.py, this is the default number of samples in the code:
self.nsamples = 2 * self.M + 2**11
Here, self.M is the number of indices where the instance is different from the background dataset.
To understand the convergence dynamics of KernelShap, we multiply this number by multiple “sample factor”s and run the algorithm. This allows us to experiment with the number of subsets sampled, and find a sample factor that ensures convergence while avoiding unnecessary extra computation. We test sample factors in the set $\{0.02,0.03,0.04,0.1,1.0,2.0,10.0\}$.
For instance, for GB1 dataset, we compute the $R^{2}$ score of the KernelShap SHAP values with different sample factors to sample factor 10 (as the ground truth).
Sample Factor Runtime $R^{2}$ w.r.t Ground Truth
0.02 4.031 0.902
0.03 4.274 0.934
0.04 4.530 0.967
0.1 5.399 0.991
1.0 19.452 0.999
2.0 34.451 1.0
10.0 160.251 1.0 ($R^{2}$ w.r.t itself trivially=1)
Table 2: Table of sample factor, runtime, and $R^{2}$ with respect to sample factor 10, on GB1 dataset.
For this dataset, the default setting, i.e., sample factor $=1$, seems good enough to make sure we are producing ground truth values and that the algorithm has converged. This is the general procedure we use for all datasets.
Appendix F Limitations & Future Work
F.1 Extending to continuous features
A limitation of FourierShap is that it currently does not natively support continuous features and they have to be handled by quantization into categorical features. However, there is value in computing efficient SHAP values for models with continuous features and this is an important potential future work. In the following, we discuss how we think such an extension would work for the two classes of models we extensively covered in this work, namely trees and MLPs.
Trees in the white-box setting. Trees can be seen as inherently discrete structures even though they perfectly work for continuous features. By setting a threshold, i.e., a continuous number to define a split of the node, all trees are inherently in fact binary. This way of thinking gives one very simple but crude extension of the current framework: To assign to each node of the tree a binary feature specifying whether a certain continuous feature is bigger or smaller than its threshold. This would increase the feature dimension of the problem to the number of nodes in the tree in the worst case scenario. With the careful design of the transformation of continuous features into node-based binary features, this could be a potential extension for which the current work can lay a foundation.
A second and more principled and less crude way to think about the case of trees is to not use a Fourier transform that we are using now which is over $Z_{2}\times Z_{2}\times...Z_{2}$ (n times where n is the number of features). But rather use a transform over $Z_{K_{1}},\times,Z_{K_{n}}$ where $K_{i}$ is the number of distinct thresholds feature $i$ has in the tree. Coming up with a closed form solution for the SHAP values in this Fourier basis is an interesting question that we also plan to look at. It is not too far-fetched to think that we can also find a closed form solution to this discrete Fourier basis. This would perfectly handle the case of any tree. Note that by a recursive formula one can readily compute the Fourier transform of trees as well.
MLPs. Regarding the case of MLPs with continuous input features, one can not simply resort to the discrete Fourier transform of type $Z_{k}\times\dots Z_{k}$ to perfectly capture this structure as it is no longer inherently discrete (like trees). This is a slightly tricky scenario because it requires us to characterize the relationship of continuous and discrete Fourier transforms of neural networks functions. In classical signal processing there are results which characterize what level of granularity in discretization is required to reconstruct continuous transform with the discrete one for a given level of error. These results depend on the Spectrum of the continuous transform. For example the case of super imposition of sinusoids, where the spectrum has a bounded domain give rise to the Nyquist rate etc.
For Neural network functions, if one had clear theoretical results on the spectral behavior of continuous transforms of neural networks it is not too hard to come up with conditions on the level of quantization required for the discrete transform to approximate them. One result that computes the spectrum appears over a multi dimensional inputs is (Theorem 1 of Rahaman et al. [2019]) but we are not aware of any spectral bias bounds on the continuous spectrum derived from this equation. However, for Neural networks with a single input dimension, bounds have been derived on the spectrum Cao et al. [2019] (in addition to the empirical results of Rahaman et al. [2019]). This intuitively means that single input Neural networks provably have a tendency to approximate functions with lower frequency sinusoids in the continuous domain. This implies that a discrete Fourier transform over $Z_{K}$ can approximate these function correctly for a sufficiently large discretization parameter K. This gives hope that a empirically a discrete Fourier transform of type $Z_{K}\times\dots Z_{K}$ for a sufficiently large $K$ would also approximate neural network function with higher dimensional inputs for a large enough discretization parameter K. Therefore, an extension of our closed-form solution for SHAP to $Z_{K}\times\dots Z_{K}$ transform could be a potential direction for the future work.
F.2 Conditional SHAP values
Another direction for future work is extending this work to compute conditional SHAP values. In this work, we only cover interventional SHAP values (the difference of interventional and conditional SHAP versions is discussed in detail in Appendix A.3).
In model-based methods like TreeSHAP Lundberg et al. [2020], one can use the tree itself to approximate conditional distributions of the data in a tractable way. One idea for extending this work is to attempt to tractably find a representation of the conditional distribution of the data using a sparse Fourier approximation of the predictor. Next, computing conditional SHAP Values using this conditional approximation.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: –
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: –
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: –
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: –
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: –
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: –
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: –
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer:[Yes]
  39. Justification: –
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: –
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: While our method may enable beneficial applications in model interpretability, it does not introduce foreseeable risks or harms.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No models or datasets with misuse risk are released.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: -
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: -
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No crowdsourcing or research with human subjects is performed in this work.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: No crowdsourcing or research with human subjects is performed in this work.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: No LLM is used as a core component in this work.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.